# Exploring Holomorphic Retracts

### v6

@@ 2. Preliminaries
We recall certain basic and general facts about retracts which would needless to say, apply to our specific cases and put it in good perspective, as well. While we would like to minimize repetition, we prefer to put the details on record to ensure a rigorous exposition. But here we repeat a few generalities from the introductory section for one last time, to cast this section systematically. Let us begin with the most general definition relevant to us: a self-map $\rho$ of a set $X$ is called a retraction if $\rho\circ\rho=\rho$; note that this idempotent condition on $\rho$ means that $\rho$ acts as the identity map on its range $Z=\rho(X)$. So, for any Hausdorff topological space $X$, its retracts, which are the images of continuous retractions, are closed subsets of $X$; henceforth Hausdorffness on $X$ will be in force. We regard $\rho$ as being non-trivial, if it is neither a constant map nor the identity. In addition to the contrasts between proper maps and retractions mentioned in the introduction, we note that non-trivial retraction mappings may very well fail to be closed maps notwithstanding their ranges being closed subsets always; indeed, as soon as $X$ is a non-compact complex manifold – the general setting of interest in this article – they always fail to be closed maps, despite the general fact that $\rho(X)$ is always closed as a subset of $X$, for every retraction map $\rho$ on $X$ as just mentioned (proofs detailed later below). However, the feature of being a quotient map is common to both retraction maps as well as proper maps. Getting back to the general case of a Hausdorff topological space $X$, if we add the assumption that $X$ is connected, then so are every of its retracts $Z$. When $X$ is furthermore contractible i.e., the space $X$ deformation retracts to a point then by composing any such deformation retraction with the retraction map $\rho$, gives a homotopy within $Z$ that reduces $Z=\rho(X)$ to a point, meaning that if $X$ is contractible then so are all of its retracts. If $X$ even happens to be a convex domain in some real topological vector space, then the retracts of $X$ are also strong deformation retracts; indeed, the homotopy $H(t,x)=(1-t)x+t\rho(x)$ substantiates this. We are ofcourse interested in holomorphic retracts: if $X$ is a complex manifold and $\rho$ is a holomorphic self-map of $X$ which is idempotent then, we say that $\rho$ is a holomorphic retraction and its image $Z=\rho(X)$ a holomorphic retract; as we shall only be talking about holomorphic retracts, we shall most often drop the adjective ‘holomorphic’. Moreover as mentioned in the introduction, our focus will be to determine the form of the retracts $Z=\rho(X)$ rather than the retractions $\rho$. A first basic and general fact about holomorphic retracts that is helpful to have in the background for the discussion in the sequel is laid down in the following well-known theorem which says that they are not just analytic varieties but actually (connected) manifolds i.e., are smooth and have the same dimension throughout.
@@
Continuing with the proof of theorem 2.1, we shall now show that in a neighbourhood of any given $p\in Z$, $Z$ is the level set of a holomorphic map whose derivative has constant rank near $p$; the implicit mapping theorem will then render the claim of the theorem. To this end then, we fix an arbitrary $p\in Z$ and passing to a holomorphic chart of the complex manifold $X$ near $p$, we may as well assume (as already mentioned), that we are in the complex Euclidean space $\mathbb{C}^{N}$, which allows us to define $\beta(x)=x-\rho(x)$ for $x$ near $p$, whose zero set is precisely $Z$ (all in the charts near $p$, to be precise). To show that the constancy of the rank of the derivative $D\beta(x)$ as mentioned, we first consider the derivative of $\rho$. The idempotence of the linear operator $D\rho(p)$ given by the above lemma, guarantees its diagonalizability as well. So, after a linear change of coordinates, $D\rho(p)$ is a diagonal matrix with diagonal entries either $1$ or 00. Let $r$ be the number of $1$’s i.e., $r={\rm rank}(D\rho(p))$. We may as well assume by a permutation of coordinates if needed, that the first principal $r\times r$ submatrix of $D\rho(p)$ is the identity matrix of order $r$, which by continuity of the determinant function implies that the determinant of the same principal minor of the nearby derivative matrices $D\rho(x)$ for all $x$ in some neighbourhood (contained in the chart) about $p$, is non-zero as well. This just means the well-known lower semi-continuity of the rank function, which when applied to $x\mapsto{\rm rank}(D\rho(x))$ implies that the rank of $D\rho(x)$ is atleast $r$ for all $x$ in some neighbourhood of $p$. Now, observe that the same arguments may be applied to $D\beta(x)$ as well. Firstly, note that $D\beta(p)=I-D\rho(p)$ where $I$ is the identity matrix of size $n$, is also an idempotent matrix, indeed diagonal with diagonal entries only $1$ or 00 and with the number of zeros being exactly $n-r$. In particular, ${\rm rank}(D\beta(p))=n-r$ and the lower semicontinuity of the rank applied to $x\mapsto{\rm rank}(D\beta(x))$ implies that the rank of $D\beta(x)$ is atleast $n-r$ in a neighbourhood of $p$. Now, we have a similar lower bound on the rank of $D\rho(x)$ as already noted, which is tempting to be translated this into an bound in-terms of $\beta$ instead of $\rho$; specifically, the upper bound ${\rm rank}(D\beta(x))\leq n-r$ in the some neighbourhood of $p$. While this is correct for $x=p$ and even along $Z$, it is not at all apriori clear as to why this bound continued to hold in an open neighbourhood of $p$. The reason for detailing this flawed argument is that it is then easy to finish the proof, within standard elementary ideas of calculus on differential manifolds. Indeed firstly, the foregoing pair of bounds when put together, only means that the rank of $D\beta(x)$ is constant in some neighbourhood of $p$, leading to the conclusion by the constant rank theorem that $Z$ is a complex manifold of dimension $r$ throughout, as we already know $Z=\rho(X)$ is connected but alas, this part of the argument is flawed, as pinpointed. This finishes the discussion of the proof of 2.1. Let us round this by noting a couple of very easy but noteworthy facts. Retracts of compact complex manifolds are compact complex submanifolds. Retracts of Stein manifolds are Stein and likewise, the property of a complex manifold being an Oka manifold, is preserved under passage to a retract thereof, as well.
@@ 2.1. Linear Retracts
[…]
@@ 2.2. Notions of extremality and strict convexity
[…]
@@
[…]
@@
For the converse, assume that $\Omega$ is $\mathbb{R}$-extremal. Let $T$ be a $\mathbb{R}-$ linear map with $\|T\|=1$. Suppose that there exist $z,w\in\overline{\Omega}$ such that $Tz=1=Tw$. Then $z,w\in\partial{\Omega}$, otherwise $T\left(z/\|z\|\right)=1/\|z\|>1$. Note that $T((1-t)z+tw)=1$. Then $(1-t)z+tw\in\partial{\Omega}$ for all $t\in[0,1]$. In particular, if $t=1/2$, we get $\frac{1}{2}(z+w)\in\partial{\Omega}$. Note then that
@@
[…]
@@
[…]
@@
Now by Claim-1, $a$ and $u$ are linearly independent. Since $span\{a\}\cap H=\{0\}$ and $span\{u\}\cap H=\{0\}$, it follows that $dim(H)<N-1$. This contradicts the fact that $H$ is a hyperplane. Hence $a+su\in\widetilde{H}$ for all $s\in\mathbb{R}$. Therefore $\|a+su\|\geq 1$. This implies that $\rm dist(a,M)\geq 1$. This completes the proof of the claim.
@@ 2.3. Geodesics and totally geodesic subsets
We recall the notion of totally geodesic subsets of a general domain (or a complex manifold) $D$ from [Kosnski]. Let $p_{1},p_{2}$ be (distinct) points in $D$; the pair $\delta=(p_{1},p_{2})$ shall be referred to as a ‘datum’. A Kobayashi extremal for $\delta$ is a holomorphic map $k:\Delta\to D$ such that both $p_{1}$ and $p_{2}$ are in the range of $k$, and the pseudo-hyperbolic distance between the preimages of $p_{1}$ and $p_{2}$ is minimized at $k$, over all holomorphic maps from $\Delta$ to $D$. We shall call the range of a Kobayashi extremal with datum $\delta$, a geodesic through $\delta$. If a set $V\subset D$ has the property that for any $p_{1},p_{2}\in V$, a geodesic through $(p_{1},p_{2})$ is contained in $V$, we shall say that $V$ is totally geodesic.
@@
Let $X$ be any complex manifold and $Z\subset X$, a retract. For a contractive pseudodistance $d$ and contractive (infinitesimal) metric $\delta$, we have
@@
It follows atleast for bounded convex domains $D\subset\mathbb{C}^{N}$ that retracts of $D$ are totally geodesic, a notion that shall be made use of, while recalling the work [Mok] of Mok. This lemma also shows that if $X$ is hyperbolic then every retract of $X$ is hyperbolic.
@@ 2.4. One dimensional Retracts
We say that a holomorphic map $\varphi:\Delta\rightarrow D$ is a complex $C$-geodesic of $D$ if for all $\zeta_{1},\zeta_{2}\in\Delta$, we have $c_{D}\left(\varphi\left(\zeta_{1}\right),\varphi\left(\zeta_{2}\right)\right)=p\left(\zeta_{1},\zeta_{2}\right)$.
@@
The following proposition relates complex geodesics with retracts.
@@
By Montel’s theorem, there exists a subsequence $f_{n_{k}}$ converging to $g$ uniformly on compact subsets of $D$. So $p\left((g\circ\varphi)(\zeta_{1}),(g\circ\varphi)(\zeta_{2})\right)=p(\zeta_{1},\zeta_{2})$ for all $\zeta_{1},\zeta_{2}\in\Delta$. By Schwarz Pick lemma, we have $g\circ\varphi\in Aut(\Delta)$. This implies that $\varphi$ has a holomorphic left inverse.
@@
This proves that $\rho$ is a holomorphic retraction map on $D$ and $\varphi(\Delta)$ is a one dimensional retract of $D$. ∎
@@
If $D$ is a simply connected Kobayashi hyperbolic domain in $\mathbb{C}^{N}$, then $1$-dimensional retracts of $D$ are precisely complex $C$-geodesics.
@@
By proposition 2.28, we get that every complex $C$-geodesic is a retract. We have only to prove that every one-dimensional retract is a complex $C$-geodesic. Let $Z$ be a $1$-dimensional retract of $D$. Note that $Z$ is a non-compact simply connected Riemann surface. By the uniformization theorem, $Z$ is biholomorphic to one of the Riemann surfaces: unit disk $\Delta,~{}\hat{\mathbb{C}},~{}\mathbb{C}$. Since $Z$ is non-compact and hyperbolic, $\hat{\mathbb{C}}$ and $\mathbb{C}$ are ruled out. Hence $Z$ is biholomorphic to $\Delta$. Let $\phi$ be a biholomorphic map from $\Delta$ to $Z$. By proposition 1.2 (the equivalence of parts a) and d) therein ) from [Dineen], we obtain that $\phi$ is a complex geodesic. So $Z$ is an image of complex $C$-geodesic. ∎
@@
Let $D$ be a balanced domain in $\mathbb{C}^{N}$. The above two definitions 2.30 and 2.31 are equivalent.
@@
Assume that $D$ is convex at $a/h(a)$ in the sense of definition 2.30. Then there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ such that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|L(a)|=h(a)$. Consider the real linear functional $:\mathbb{C}^{N}\to\mathbb{R}$ defined by $T(z)=Re(\alpha L(z))$, where $\alpha=|L(a)|/L(a)$. Note that $|\alpha|=1$ and we obtain that
@@
Conversely, assume that $D$ is convex at $a/h(a)$ as per definition 2.31. Then there exists a $\mathbb{R}$-linear functional $f:\mathbb{C}^{N}\to\mathbb{R}$ with $|f(z)|\leq h(z)$ and $|f(a)|=h(a)$. Define a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ by $L(z)=f(z)-if(iz)$. Observe that for each $z\in\mathbb{C}^{N}$, there exist $\lambda_{z}\in\mathbb{C}$ with $|\lambda_{z}|=1$ such that $\lambda_{z}L(z)=|L(z)|$. Note that

### v7

@@ 2. Preliminaries
We recall certain basic and general facts about retracts which would needless to say, apply to our specific cases and put it in good perspective, as well. While we would like to minimize repetition, we prefer to put the details on record to ensure a rigorous exposition. We believe many results here (if not all) are probably folklore or even if new, are relatively straightforward but essential. However, we repeat here a few generalities from the introductory section for one last time, to cast this section and the rest, systematically. Let us begin with the most general definition relevant to us: a self-map $\rho$ of a set $X$ is called a retraction if $\rho\circ\rho=\rho$; note that this idempotent condition on $\rho$ means that $\rho$ acts as the identity map on its range $Z=\rho(X)$. So, for any Hausdorff topological space $X$, its retracts, which are the images of continuous retractions, are closed subsets of $X$; henceforth Hausdorffness on $X$ will be in force. We regard $\rho$ as being non-trivial, if it is neither a constant map nor the identity. In addition to the contrasts between proper maps and retractions mentioned in the introduction, we note that non-trivial retraction mappings may very well fail to be closed maps notwithstanding their ranges being closed subsets always; indeed, as soon as $X$ is a non-compact complex manifold – the general setting of interest in this article – they always fail to be closed maps, despite the general fact that $\rho(X)$ is always closed as a subset of $X$, for every retraction map $\rho$ on $X$ as just mentioned (proofs detailed later below). However, the feature of being a quotient map is common to both retraction maps as well as proper maps. Getting back to the general case of a Hausdorff topological space $X$, if we add the assumption that $X$ is connected, then so are every of its retracts $Z$. When $X$ is furthermore contractible i.e., the space $X$ deformation retracts to a point then by composing any such deformation retraction with the retraction map $\rho$, gives a homotopy within $Z$ that reduces $Z=\rho(X)$ to a point, meaning that if $X$ is contractible then so are all of its retracts. If $X$ even happens to be a convex domain in some real topological vector space, then the retracts of $X$ are also strong deformation retracts; indeed, the homotopy $H(t,x)=(1-t)x+t\rho(x)$ substantiates this. We are ofcourse interested in holomorphic retracts: if $X$ is a complex manifold and $\rho$ is a holomorphic self-map of $X$ which is idempotent then, we say that $\rho$ is a holomorphic retraction and its image $Z=\rho(X)$ a holomorphic retract; as we shall only be talking about holomorphic retracts, we shall most often drop the adjective ‘holomorphic’. Moreover as mentioned in the introduction, our focus will be to determine the form of the retracts $Z=\rho(X)$ rather than the retractions $\rho$. A first basic and general fact about holomorphic retracts that is helpful to have in the background for the discussion in the sequel is laid down in the following well-known theorem which says that they are not just analytic varieties but actually (connected) manifolds i.e., are smooth and have the same dimension throughout.
@@
Continuing with the proof of theorem 2.1, we shall now show that in a neighbourhood of any given $p\in Z$, $Z$ is the level set of a holomorphic map whose derivative has constant rank near $p$; the implicit mapping theorem will then render the claim of the theorem. To this end then, we fix an arbitrary $p\in Z$ and passing to a holomorphic chart of the complex manifold $X$ near $p$, we may as well assume (as already mentioned), that we are in the complex Euclidean space $\mathbb{C}^{N}$, which allows us to define $\beta(x)=x-\rho(x)$ for $x$ near $p$, whose zero set is precisely $Z$ (all in the charts near $p$, to be precise). To show that the constancy of the rank of the derivative $D\beta(x)$ as mentioned, we first consider the derivative of $\rho$. The idempotence of the linear operator $D\rho(p)$ given by the above lemma, guarantees its diagonalizability as well. So, after a linear change of coordinates, $D\rho(p)$ is a diagonal matrix with diagonal entries either $1$ or 0. Let $r$ be the number of $1$’s i.e., $r={\rm rank}(D\rho(p))$. We may as well assume by a permutation of coordinates if needed, that the first principal $r\times r$ submatrix of $D\rho(p)$ is the identity matrix of order $r$, which by continuity of the determinant function implies that the determinant of the same principal minor of the nearby derivative matrices $D\rho(x)$ for all $x$ in some neighbourhood (contained in the chart) about $p$, is non-zero as well. This just means the well-known lower semi-continuity of the rank function, which when applied to $x\mapsto{\rm rank}(D\rho(x))$ implies that the rank of $D\rho(x)$ is atleast $r$ for all $x$ in some neighbourhood of $p$. Now, observe that the same arguments may be applied to $D\beta(x)$ as well. Firstly, note that $D\beta(p)=I-D\rho(p)$ where $I$ is the identity matrix of size $n$, is also an idempotent matrix, indeed diagonal with diagonal entries only $1$ or 0 and with the number of zeros being exactly nrn-r. In particular, ${\rm rank}(D\beta(p))=n-r$ and the lower semicontinuity of the rank applied to $x\mapsto{\rm rank}(D\beta(x))$ implies that the rank of $D\beta(x)$ is atleast nrn-r in a neighbourhood of $p$. Now, we have a similar lower bound on the rank of $D\rho(x)$ as already noted, which is tempting to be translated this into an bound in-terms of $\beta$ instead of $\rho$; specifically, the upper bound ${\rm rank}(D\beta(x))\leq n-r$ in the some neighbourhood of $p$. While this is correct for $x=p$ and even along $Z$, it is not at all apriori clear as to why this bound continued to hold in an open neighbourhood of $p$. The reason for detailing this flawed argument is that it is then easy to finish the proof, within standard elementary ideas of calculus on differential manifolds. Indeed firstly, the foregoing pair of bounds when put together, only means that the rank of $D\beta(x)$ is constant in some neighbourhood of $p$, leading to the conclusion by the constant rank theorem that $Z$ is a complex manifold of dimension $r$ throughout, as we already know $Z=\rho(X)$ is connected but alas, this part of the argument is flawed, as pinpointed. This finishes the discussion of the proof of 2.1. Let us round this by noting a couple of very easy but noteworthy facts. Retracts of compact complex manifolds are compact complex submanifolds. Retracts of Stein manifolds are Stein and likewise, the property of a complex manifold being an Oka manifold, is preserved under passage to a retract thereof, as well. In addition to the contrasts between retractions and proper mappings mentioned in the introduction, we have the following.
@@
Since for a good part of the article, we deal with circular domains, we state one of the most famous results about mappings between such domains which will be used in numerous occasions in this article; we shall simply refer to this theorem due to H.Cartan as Cartan’s theorem.
Theorem 2.6.
(H. Cartan) Suppose $\Omega_{1},\Omega_{2}$ are circular domains in $\mathbb{C}^{N}$ with $0\in\Omega_{1}\cap\Omega_{2}$ and $\Omega_{1}$ is bounded. If $F$ is a biholomorphic map of $\Omega_{1}$ onto $\Omega_{2}$ with $F(0)=0$, then $F$ is a linear transformation.
There shall also be occasions where another result of H.Cartan will be used, as stated next. We shall refer to it as the classical uniqueness theorem of Cartan.
Theorem 2.7.
Suppose $\Omega$ is a bounded domain in $\mathbb{C}^{N}$ and $F:\Omega\to\Omega$ is a holomorphic map. If for some $p\in\Omega,~F(p)=p$ and $F^{\prime}(p)=I$, then $F(z)=z$ for all $z\in\Omega$.
2.1. Balanced Domains in $\mathbb{C}^{N}$
In this subsection, we recall the basic properties of balanced domain and its associated Minkowski function, following [Jrncki_invrnt_dst].
Definition 2.8.
Let $D\subset\mathbb{C}^{N}$ be a balanced domain. Define its Minkowski function $h_{D}:\mathbb{C}^{N}\longrightarrow\mathbb{R}^{+}$,
$h_{D}(z):=\inf\{t>0:z/t\in D\},\quad z\in\mathbb{C}^{N}.$
Proposition 2.9.
We have with the notations as in the above definition, the following properties of the Minkowski functional.
  1. (a)
    $h_{D}(\lambda z)=|\lambda|h_{D}(z)$, for all $z\in\mathbb{C}^{N},\lambda\in\mathbb{C}$.
  2. (b)
    $D=\left\{z\in\mathbb{C}^{N}:h_{D}(z)<1\right\}$.
  3. (c)
    If $h:\mathbb{C}^{N}\longrightarrow\mathbb{R}^{+}$is an upper semicontinuous function such that $h(\lambda z)=|\lambda|h(z),z\in\mathbb{C}^{N},\lambda\in\mathbb{C}$, then the set $D:=\left\{z\in\mathbb{C}^{N}:h(z)<1\right\}$ is a balanced domain and $h\equiv h_{D}$.
  4. (d)
    If $h_{D}$ is continuous, then $\partial D=\left\{z\in\mathbb{C}^{N}:h_{D}(z)=1\right\}$.
  5. (e)
    $D$ is convex if and only if $h_{D}$ is a seminorm, i.e., $h_{D}(z+w)\leq h_{D}(z)+h_{D}(w),z,w\in\mathbb{C}^{N}$.
  6. (f)
    If $D$ is bounded, then $h_{D}^{-1}(0)=\{0\}$.
An important little observation that will be particularly useful to us is the following. Fix any norm $\|\cdot\|$ on $\mathbb{C}^{N}$. Then, for any $z\in\mathbb{C}^{N}$, the Minkowski functional of $D$ can be expressed as:
$h_{D}(z)\;=\;\frac{\|z\|}{\|\pi(z)\|},$
where $\pi(z)=\alpha_{0}z\in\partial D$ with $\alpha_{0}:=\sup\{\alpha>0:\alpha z\in D\}$.
@@ 2.2. Linear Retracts
It suffices to assume that $D_{L_{1}\oplus L_{2}}$ is a linear retract of $D$ for every pair of (complex) lines $L_{1},L_{2}$ through the origin $\mathbb{C}^{N}$ to derive a contradiction. Indeed with such an assumption, we would have that every $2$ dimensional subspace of $D$ is a linear retract of $D$. But then, this would imply by theorem-A in [Bhnblst] that $(\mathbb{C}^{N},\|\cdot\|)$ is a Hilbert space, where $\|\cdot\|$ denotes the Minkowski functional on $D$ which in turn implies the existence of a unitary map $T$ from $(\mathbb{C}^{N},\|\cdot\|)$ to $(\mathbb{C}^{N},\|\cdot\|_{\ell^{2}})$. Therefore $T(D)=\mathbb{B}$, contradicting the hypothesis that $D$ is not equivalent to $\mathbb{B}$. ∎
@@ 2.3. Notions of extremality and strict convexity
Let $\Omega$ be a balanced bounded convex domain in $\mathbb{C}^{N}$; so, $\mathbb{C}^{N}$ may be regarded as a Banach space with respect to the norm $\|\cdot\|$ whose unit ball is $\Omega$. We say that $\Omega$ is strictly $K$-convex if for every $K$-linear functional $T$ with $\|T\|=1$, corresponds just one $z\in\overline{\Omega}$ such that $Tz=1$.
@@
Let $\Omega$ be a balanced bounded convex domain in $\mathbb{C}^{N}$; so, $\mathbb{C}^{N}$ may be regarded as a Banach space with respect to the norm $\|\cdot\|$ whose unit ball is $\Omega$. We say that $\Omega$ is rotund if $tx_{1}+(1-t)x_{2}\in\Omega$ whenever $x_{1}$ and $x_{2}$ are different points of $\partial\Omega$ and $0<t<1$.
@@
For the converse, assume that $\Omega$ is $\mathbb{R}$-extremal. Let $T$ be a $\mathbb{R}$- linear map with $\|T\|=1$. Suppose that there exist $z,w\in\overline{\Omega}$ such that $Tz=1=Tw$. Then $z,w\in\partial{\Omega}$, otherwise $T\left(z/\|z\|\right)=1/\|z\|>1$. Note that T((1t)z+tw)=1T((1-t)z+tw)=1. Then $(1-t)z+tw\in\partial{\Omega}$ for all $t\in[0,1]$. In particular, if $t=1/2$, we get $\frac{1}{2}(z+w)\in\partial{\Omega}$. Note then that
@@
Let $\Omega$ be a bounded balanced convex domain in $\mathbb{C}^{N}$. Note that $\Omega$ is the unit ball with respect to some norm $\|\cdot\|$ in $\mathbb{C}^{N}$. Assume $\Omega$ is strictly $\mathbb{R}$-convex. We want to prove that $\Omega$ is strictly $\mathbb{C}$-convex. Let $T$ be $\mathbb{C}$ -linear functional with $\|T\|=1$ such that $T(z)=T(w)=1$ for some $z,w\in\overline{\Omega}$. Every $\mathbb{C}$ -linear functional must be an $\mathbb{R}$-linear functional. Then $T$ is a $\mathbb{R}$-linear functional with $\|T\|=1$ and $T(z)=T(w)=1$. Since $\Omega$ is strictly $\mathbb{R}$-convex, we get $z=w$.
@@
Claim-2: ${\rm dist}(a,M)=1$, where $\rm dist$} denotes the distance induced by the norm $\|\cdot\|$ with respect to which $\Omega$ is the unit ball.
@@
Now by Claim-1, $a$ and $u$ are linearly independent. Since $span\{a\}\cap H=\{0\}$ and $span\{u\}\cap H=\{0\}$, it follows that dim(H)<N1dim(H)<N-1. This contradicts the fact that $H$ is a hyperplane. Hence $a+su\in\widetilde{H}$ for all $s\in\mathbb{R}$. Therefore $\|a+su\|\geq 1$. This implies that $\rm dist(a,M)\geq 1$. This completes the proof of the claim.
@@ 2.4. Geodesics and totally geodesic subsets
We recall the notion of totally $K$-geodesic subsets of a general domain (or a complex manifold) $D$ from [Kosnski]. Let $p_{1},p_{2}$ be (distinct) points in $D$; the pair $\delta=(p_{1},p_{2})$ shall be referred to as a ‘datum’. A Kobayashi extremal for $\delta$ is a holomorphic map $k:\Delta\to D$ such that both $p_{1}$ and $p_{2}$ are in the range of $k$, and the pseudo-hyperbolic distance (also known as the Möbius distance on $\Delta$) between the preimages of $p_{1}$ and $p_{2}$ is minimized at $k$, over all holomorphic maps from $\Delta$ to $D$. We shall call the range of a Kobayashi extremal with datum $\delta$, a $K$-geodesic through $\delta$. If a set $V\subset D$ has the property that for any $p_{1},p_{2}\in V$, a $K$-geodesic through $(p_{1},p_{2})$ is contained in $V$, we shall say that $V$ is totally $K$-geodesic. A similar definition can be made with invariant metrics or distances, in particular the Caratheodory distance, but we do not pursue such matter here.
@@
Let $X$ be any complex manifold and $Z\subset X$, a retract. For a contractive pseudo-distance $d$ and contractive (infinitesimal) pseudo-metric $\delta$, we have
@@
It follows atleast for bounded convex domains $D\subset\mathbb{C}^{N}$ that retracts of $D$ are totally geodesic; the more standard definition of ‘totally geodesic’ shall also be made use of, while recalling the work [Mok] of Mok. Getting back to the above lemma, it also shows that if $X$ is hyperbolic then every retract of $X$ is hyperbolic.
@@ 2.5. One dimensional Retracts
We say that a holomorphic map $\varphi:\Delta\rightarrow D$ is a complex $C$-geodesic of $D$ if for all $\zeta_{1},\zeta_{2}\in\Delta$, we have $c_{D}\left(\varphi\left(\zeta_{1}\right),\varphi\left(\zeta_{2}\right)\right)=p\left(\zeta_{1},\zeta_{2}\right)$. Occasionally, by abuse of language, we shall also refer to the image of $\varphi$ as a complex $C$-geodesic.
@@
The following proposition relates complex $C$-geodesics with retracts. This seems well-known; it can be found stated in [Dineen]. However, the proof details seem not to have been recorded anywhere which we therefore include here; particularly because it is a basic result about one-dimensional retracts.
@@
By Montel’s theorem, there exists a subsequence $f_{n_{k}}$ converging to $g$ uniformly on compact subsets of $D$. So $p\left((g\circ\varphi)(\zeta_{1}),(g\circ\varphi)(\zeta_{2})\right)=p(\zeta_{1},\zeta_{2})$ for all $\zeta_{1},\zeta_{2}\in\Delta$. By Schwarz Pick lemma, we have $g\circ\varphi\in{\rm Aut}(\Delta)$. This implies that $\varphi$ has a holomorphic left inverse.
@@
This proves that $\rho$ is a holomorphic retraction map on $D$ and $\varphi(\Delta)$ is a one-dimensional retract of $D$. ∎
It may be surprising that the hypothesis has no condition about $D$. Indeed, it may happen that $c_{D}\equiv 0$, there are no complex $C$- geodesic and the result is vacuously true. The converse is true for the class of domains of much interest in this article namely, bounded balanced pseudoconvex domains. More generally, we have the following.
@@
If $D$ is a simply connected Kobayashi hyperbolic domain in $\mathbb{C}^{N}$, then the one-dimensional retracts of $D$ are precisely its complex $C$-geodesics.
@@
By proposition 2.32, we get that every complex $C$-geodesic is a retract. We have only to prove that every one-dimensional retract is a complex $C$-geodesic. Let $Z$ be a $1$-dimensional retract of $D$. Note that $Z$ is a non-compact simply connected Riemann surface. By the uniformization theorem, $Z$ is biholomorphic to one of the Riemann surfaces: unit disk $\Delta,~\hat{\mathbb{C}}$$\mathbb{C}$. Since $Z$ is non-compact and hyperbolic, $\hat{\mathbb{C}}$ and $\mathbb{C}$ are ruled out. Hence $Z$ is biholomorphic to $\Delta$. Let $\phi$ be a biholomorphic map from $\Delta$ to $Z$. By proposition 1.2 (the equivalence of parts a) and d) therein ) from [Dineen] (proposition 11.1.7 in [Jrncki_invrnt_dst]), we obtain that $\phi$ is a complex $C$-geodesic. So $Z$ is an image of a complex $C$-geodesic. ∎
@@
Let $D$ be a balanced domain in $\mathbb{C}^{N}$. The above two definitions 2.34 and 2.35 are equivalent.
@@
Assume that $D$ is convex at $a/h(a)$ in the sense of definition 2.34. Then there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ such that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|L(a)|=h(a)$. Consider the real linear functional $\mathbb{C}^{N}\to\mathbb{R}$ defined by $T(z)=Re(\alpha L(z))$, where $\alpha=|L(a)|/L(a)$. Note that $|\alpha|=1$ and we obtain that
@@
Conversely, assume that $D$ is convex at $a/h(a)$ as per definition 2.35. Then there exists a $\mathbb{R}$-linear functional $f:\mathbb{C}^{N}\to\mathbb{R}$ with $|f(z)|\leq h(z)$ and $|f(a)|=h(a)$. Define a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ by L(z)=f(z)if(iz)L(z)=f(z)-if(iz). Observe that for each $z\in\mathbb{C}^{N}$, there exist $\lambda_{z}\in\mathbb{C}$ with $|\lambda_{z}|=1$ such that $\lambda_{z}L(z)=|L(z)|$. Note that